CO_{2}-BROADENED H_{2}O IN THE PURE ROTATION AND ν_{2} FUNDAMENTAL REGIONS.

\underline{LINDA R. BROWN}^{a}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; CATLIN M. HUMPHREY and ROBERT R. GAMACHE, Department of Environmental, Earth & Atmospheric Sciences, University of Massachusetts School of Marine Sciences, 265 Riverside Street, Lowell, MA 01854-5045.

The CO_{2}-broadened water coefficients (half-widths, line shifts and temperature dependence of the widths) are predicted using a fully complex Robert-Bonamy formulation for the 937 allowed and forbidden perpendicular type transitions of (000)-(000) between 200-900 cm^{−1} in order to facilitate atmospheric remote sensing of the Mars and Venus. In addition, empirical Lorentz linewidths and pressure-induced frequency-shifts of CO_{2}-broadened H_{2}^{16}O are obtained at room temperature for 257 perpendicular transitions of the (010)-(000) fundamental. For this, calibrated spectra recorded at 0.0054 cm^{−1} resolution are measured assuming Voigt line shapes. For transitions between 1287 and 1988 cm^{−1} with rotational quanta up to J = 13 and Ka = 6, the widths vary from 0.045 to 0.212 cm^{−1} atm^{−1} at 300 K; the pressure-shifts are quite large and range from -0.0386 to +0.0436 cm^{−1} atm^{−1}. Observed and calculated values are compared.

aSupport by NASA and NSF is gratefully acknowledged